# The functional Loomis-Whitney type inequality in the Heisenberg groups and Projection theorems over finite fields

### v1

@@ Abstract
We develop a functional Loomis-Whitney framework on the finite Heisenberg groups $\mathbb{H}^{n}(\mathbb{F}_{q})$ and discover connections to the boundedness and orthogonal projection problems. For $n=1$ we determine the sharp region of exponents $(u_{1},u_{2})$ for which the associated bilinear projection form is bounded uniformly in $q$, namely
@@
which includes the endpoint $L^{3/2}\times L^{3/2}\to L^{1}$. For general $n$ we prove a multilinear estimate at the critical exponent
$u=\frac{n(2n+1)}{n+1},$
via an induction on $n$ that exploits the group’s fiber structure together with multilinear interpolation. Specializing to indicators yields a sharp Loomis-Whitney type set inequality that controls $|K|$ for every finite $K\subset\mathbb{H}^{n}(\mathbb{F}_{q})$ by the sizes of its $2n$ Heisenberg projections $\{\pi_{j}(K)\}$, forcing a large projection in every configuration. A straightening map then converts these bounds into covering statements by additive cosets and provides a new approach to the orthogonal projection problem onto the vertical hyperplanes $\{x_{j}=0\}$, which presents an interesting link between commutative and non-commutative settings. The obtained results are optimal up to absolute constants, and in the planar case $n=1$, when the size of the set is not too small, our bound can be sharpened further using a point-line incidence estimate.
@@ 1 Introduction
The Loomis–Whitney inequality ([18]) bounds the volume of a $d$-dimensional body in $\mathbb{R}^{d}$ by the volumes of its (d1)(d-1)-dimensional coordinate projections. In the discrete setting it reads as follows. For $d\geq 2$, let $B_{1},\dots,B_{d}$ be finite sets, and let $A\subseteq B_{1}\times\cdots\times B_{d}$. For $i\in[d]$, define
@@
There are powerful entropy-theoretic generalizations of the inequality (1), notably Han’s inequality and Shearer’s inequality [19, 11]. Variants of this inequality, in relation to the set addition, also appear in additive combinatorics for abelian groups with numerous applications [21, 10].
This paper investigates such inequalities in the noncommutative setting, focusing on the Heisenberg group over a finite field. In this context, such inequalities are not only of independent interest but also shed light on group structures and related combinatorial problems. Some earlier works on the Heisenberg group can be found in [13, 14, 20, 15, 1]. Interestingly, this topic is connected to questions on boundedness and orthogonal projection problems. The approach developed here offers a new perspective on orthogonal projections onto coordinate planes. Further details will be discussed at the end of this section.
In the continuous setting, the Loomis–Whitney–type inequalities for the Heisenberg ( and more generally Carnot) groups have been studied extensively with applications in multilinear restriction theory and related areas; for example, see [8, 24, 3] and references therein.
@@
By definition, the group $\mathbb{H}^{n}(\mathbb{F}_{q})$ and the vector space $\mathbb{F}_{q}^{2n+1}$ have the same underlying sets. Thus, whenever necessary, we shall identify them only as sets: $\mathbb{H}^{n}(\mathbb{F}_{q})=\mathbb{F}_{q}^{2n+1}$. For each $1\leq j\leq 2n$, let
@@
be the coordinate line corresponding to $x_{j}$, where $e_{j}$ ($j=1,\ldots,2n$) form the standard basis of $\mathbb{F}_{q}^{2n+1}.$
Then we easily see that $W_{j}$ and $L_{j}$ are subgroups of $\mathbb{H}^{n}(\mathbb{F}_{q})$ for $j=1,\ldots,2n$. Further, $L_{j}$ is the orthogonal complement $W_{j}^{\perp}$ of $W_{j}$, w.r.t the standard inner product on $\mathbb{F}_{q}^{2n+1}$; see Subsection 2.2 for these materials. The subgroups $W_{j}$ and $L_{j}$ will be often referred to as vertical and horizontal, respectively. We will provide more details on subgroup structures of $\mathbb{H}^{n}(\mathbb{F}_{q})$ in Section 2.
For $x\in\mathbb{F}^{2n}_{q}$ and $j\in\{1,\ldots,2n\}$ we use the notation $\widehat{x}_{j}$ to represent the element in $\mathbb{F}^{2n-1}_{q}$ obtained by entirely removing the $j$-th coordinate of $x$ while the notation $\check{x}_{j}$ to represent the element in $\mathbb{F}^{2n}_{q}$ obtained by setting the $j$-th coordinate of $x$ to 0. Let
@@
be the projections, respectively, defined by
@@
A direct computation shows that for each $j=1,\ldots,2n$, an element $(x,t)$ of $\mathbb{H}^{n}(\mathbb{F}_{q})$ can be written
@@
The fiber of $\pi_{j}$ at a point $(u,\tau)$ in $W_{j}$ is a line in the affine space $\mathbb{F}_{q}^{2n+1}$ passing through $(u,\tau)$. More precisely, for $j=1,\ldots,n$, the fiber of $\pi_{j}$ at $(u,\tau)\in W_{j}$ (so $u_{j}=0$) is given by
@@
That is, for each $j=1,\ldots,2n,$ the fibers of $\pi_{j}$ are exactly cosets of $W_{j}^{\perp}=L_{j}$.
@@
Throughout, for each $j=1,\ldots,2n,$ we shall identify $W_{j}=\mathbb{F}_{q}^{2n}$ via the map $\phi_{j}:W_{j}\stackrel{{\scriptstyle\sim}}{{\rightarrow}}\mathbb{F}_{q}^{2n}$ defined by $(\check{x}_{j},t)\mapsto(\widehat{x}_{j},t)$. For simplicity, for a function $f$ on $\mathbb{F}_{q}^{2n}$ and $(x,t)\in\mathbb{H}^{n}(\mathbb{F}_{q})$, we write $f(\pi_{j}(x,t))$ for $f(\phi_{j}(\pi_{j}(x,t)))$.
@@
Let $n$ be a positive integer. For extended real numbers $1\leq u_{1},u_{2},\ldots,u_{2n}\leq\infty$, define $LW(u_{1},u_{2},\ldots,u_{2n})$ as the smallest constant, if it exists, such that the inequality
@@
holds for all nonnegative $f_{1},f_{2},\ldots,f_{2n}\in\mathscr{F}(\mathbb{F}_{q}^{2n})$. If it does not exist, then we set $LW(u_{1},u_{2},\ldots,u_{2n}):=\infty.$
When $f_{1}=\cdots=f_{2n}$ are all the indicator functions of a subset $K$ of $\mathbb{H}^{n}(\mathbb{F}_{q})$, then the above inequality presents the Loomis-Whitney type inequality.
@@
In this article we write $X\lesssim Y$ to mean that there exists some constant $C>0$ such that $X\leq CY$.
@@
For any nonnegative $f_{1},f_{2}\in\mathscr{F}(\mathbb{F}_{q}^{2})$, we have
$\frac{1}{q^{3}}\sum_{(x,t)\in\mathbb{H}^{1}(\mathbb{F}_{q})}f_{1}(\pi_{1}(x,t))f_{2}(\pi_{2}(x,t))\lesssim\|f_{1}\|_{L^{3/2}(\mathbb{F}_{q}^{2},dx)}\|f_{2}\|_{L^{3/2}(\mathbb{F}_{q}^{2},dx)}.$
When $n>1$, it is difficult to find necessary and sufficient conditions on $u_{1},u_{2},\ldots,u_{2n}$ under which $LW(u_{1},u_{2},\ldots,u_{2n})\lesssim 1$, due to the lack of understanding of structures of the group $\mathbb{H}^{n}(\mathbb{F}_{q})$. However, by using Corollary 1.3 and the inductive argument introduced in the continuous case [8], we are able to prove the following generalization of Corollary 1.3 for $u_{1}=\cdots=u_{2n}=\frac{n(2n+1)}{n+1}$.
@@
Without making additional assumptions, we can easily check that the upper bound for $|K|$ in (9) is sharp. Indeed, for $n=1,$ fix $t_{0}\in\mathbb{F}_{q}$ and let $K:=\{(x_{1},0,t_{0})\colon x_{1}\in\mathbb{F}_{q}\}$. Then we have $|K|=q$. Moreover, $|\pi_{1}(K)|=1$ and $|\pi_{2}(K)|=q$. Thus, the bound $|K|\leq q^{\frac{1}{3}}|\pi_{1}(K)|^{\frac{2}{3}}|\pi_{2}(K)|^{\frac{2}{3}}$ is sharp.
In light of this example, we cannot replace the factor $q^{\frac{1}{2n+1}}$ by $|\pi_{v}(K)|^{\frac{1}{2n+1}}$, where $\pi_{v}(K)$ is the image of $K$ under the (vertical) projection $\pi_{v}:\mathbb{H}^{n}(\mathbb{F}_{q})\rightarrow\mathbb{F}_{q}$ defined by $(x,t)\mapsto t$. This example can be extended to larger $n$ by setting
$K\;=\;\Big\{(x,t_{0})\in\mathbb{H}^{n}(\mathbb{F}_{q}):\ x_{n+1}=\cdots=x_{2n}=0\Big\}.$
Then $|K|=q^{n}$, and for the vertical projections $\pi_{1},\dots,\pi_{2n}$ we have
Consequently,
$q^{n}\;=\;|K|=q^{\frac{1}{2n+1}}\prod_{j=1}^{2n}|\pi_{j}(K)|^{\frac{n+1}{n(2n+1)}}.$
When $n=1$ and $|K|>q$, by using the point-line incidence theorem due to Vinh [23], we are able to provide a better lower bound for the quantity $\max\{|\pi_{1}(K)|,|\pi_{2}(K)|\}$.
@@
$\frac{|\pi_{1}(K)|\,|\pi_{2}(K)|}{q}\;+\;q^{1/2}\sqrt{|\pi_{1}(K)|\,|\pi_{2}(K)|}\;=\;m^{2}q\;+\;mq^{3/2},$
@@
Now we address how our results are connected to some questions in the literature.
Bhowmik, Iosevich, Koh, and Pham’s boundedness problem:
Let $\mathcal{K}\colon\mathbb{F}_{q}^{d}\times\mathbb{F}_{q}^{d}\to\mathbb{R}$ be any function (which, below, will play a role of the kernel of an integral operator for the counting measure), and $\mathcal{G}=(V,E)$ a connected ordered graph on $n$ vertices. In a recent paper [4], Bhowmik, Iosevich, Koh, and Pham introduced the boundedness problem for a multilinear operator associated with $\mathcal{K}$ and $\mathcal{G}$. To be precise, define
by
The operator $\Lambda_{\mathcal{G}}^{\mathcal{K}}$ is a general form of several geometric counting functions:
  1. 1.
  2. 2.
  3. 3.
    If $\mathcal{G}$ is a tree and $f_{1}=\cdots=f_{m}$ are the indicator functions of a set $A$ in $\mathbb{F}_{q}^{d}$, then the operator $\Lambda_{\mathcal{G}}^{\mathcal{K}}$ counts the number of copies of the given tree $\mathcal{G}$ in $A$.
The inequality (7) for $n=1$ is essentially the same as the inequality (10) for $m=2$, with suitable $\mathcal{G}$ and $\mathcal{K}$ chosen. To see this, define $\phi\colon\mathbb{F}_{q}^{3}\to\mathbb{F}_{q}^{3}$ by $\phi(x,y,t)=(x,y,t+\frac{xy}{2})$. Then, since $\phi$ is a bijective map, the left hand side of (7) can be written
Then the left hand side $\Lambda_{\mathcal{G}}^{\mathcal{K}}(f_{1},f_{2})$ of the inequality (10) is equal to
which coincides with the value (1). Thus, in this case, both inequalities (7) and (10) coincide with each other. This connection suggests a further study of the operator $\Lambda_{\mathcal{G}}^{\mathcal{K}}$ for various $\mathcal{K}$ and $\mathcal{G}$.
The orthogonal projection problem:
Fix a subset $K\subset\mathbb{F}_{q}^{2n+1}$. For $r\geq 1,$ let
[…]
where, by translates of $W^{\perp}$, we mean $\{x+W^{\perp}\colon x\in\mathbb{F}_{q}^{2n+1}\}$. Chen [6] proved that
(Informally, it states that for most subspaces $W\subset\mathbb{F}_{q}^{2n+1}$, the set $K$ cannot be covered by a relatively “small” number of translates of $W^{\perp}.$)
Let $\mathscr{H}=\{W_{1},\dots,W_{2n}\}$ be the set of the vertical hyperplanes of $\mathbb{H}^{n}(\mathbb{F}_{q})$, given in (3). If $|\mathscr{H}|>|\mathscr{E}_{r}|$, then $\mathscr{H}\nsubseteq\mathscr{E}_{r}$ and hence there exists $W_{j}\in\mathscr{H}$ such that $K$ cannot be covered by $\leq r$ translates of $W_{j}^{\perp}$.
Define $T_{j}:\mathbb{F}_{q}^{2n+1}\to\mathbb{F}_{q}^{2n+1}$ by
This map is well-defined and bijective. We observe that for every $j\in\{1,\dots,2n\}$ and every $(u,\tau)\in W_{j}$,
$T_{j}\big(\pi_{j}^{-1}\{(u,\tau)\}\big)=(u,\tau)+W_{j}^{\perp},$
i.e. $T_{j}$ maps each Heisenberg fiber bijectively onto an additive translate of $W_{j}^{\perp}$.
Notice that $|T_{j}(K)|=|K|$. Thus, the minimal number of translates of $W_{j}^{\perp}$ that cover $T_{j}(K)$ equals $|\pi_{j}(K)|$. Therefore
$|\mathscr{H}|>|\mathscr{E}_{r}|\quad\Longrightarrow\quad\exists\,j\in\{1,\dots,2n\}:\;|\pi_{j}(K)|>r.$ (14)
To describe the connection between some of our results and Chen’s, assume that
$r\;<\;\frac{n\,|K|}{q^{2n}}.$ (15)
Then from the second line of (12), we easily obtain $|\mathscr{E}_{r}|<2n=|\mathscr{H}|$. Thus, by (14) there exists some $j$ with $|\pi_{j}(K)|>r$.
Since this holds for every $r$ satisfying (15), we obtain
$\quad\max_{1\leq j\leq 2n}|\pi_{j}(K)|\;\gtrsim_{n}\;\frac{|K|}{q^{2n}}\quad$
(up to a constant depending only on $n$). This lower bound extracted from Chen’s result for the tiny family $\mathscr{H}$, is substantially weaker than our Heisenberg Loomis–Whitney bound $\max_{j}|\pi_{j}(K)|\;\gtrsim\;|K|^{\frac{2n+1}{2(n+1)}}\,q^{-\frac{1}{2(n+1)}},$ given in Corollary 1.5. To be precise,
$\frac{|K|^{\frac{2n+1}{2(n+1)}}q^{-\frac{1}{2(n+1)}}}{|K|/q^{2n}}\;=\;|K|^{-\frac{1}{2n+2}}\;q^{\,2n-\frac{1}{2(n+1)}}\;\geq\;\,q^{\,2n-1}\,.$
Since $T_{j}$ are bijective maps, the method developed in this paper offers a new approach to address the orthogonal projection problem on the family of vertical hyperplanes.
In light of these observations, it is natural to investigate projection problems in the Heisenberg group, paralleling the continuous theory of Balogh, Fassler, Mattila, and Tyson [2]. We will develop this in a sequel paper.

### v2

# A functional Loomis–Whitney type inequality in the Heisenberg group and projection theorems over finite fields

@@ Abstract
We establish functional Loomis–Whitney type inequalities in the finite Heisenberg group $\mathbb{H}^{n}(\mathbb{F}_{q})$. For $n=1$, we determine the sharp region of exponents $(u_{1},u_{2})$ for which the Heisenberg Loomis–Whitney inequality
$\frac{1}{q^{3}}\sum_{(x,t)\in\mathbb{H}^{1}(\mathbb{F}_{q})}f_{1}(\pi_{1}(x,t))\,f_{2}(\pi_{2}(x,t))\;\lesssim\;\|f_{1}\|_{L^{u_{1}}(\mathbb{F}_{q}^{2},dx)}\|f_{2}\|_{L^{u_{2}}(\mathbb{F}_{q}^{2},dx)}$
holds uniformly in $q$, namely
@@
which includes the endpoint estimate $L^{\frac{3}{2}}\times L^{\frac{3}{2}}\to L^{1}$. For general $n$, we prove the symmetric multilinear estimate at the endpoint exponent $u=\frac{n(2n+1)}{n+1},$ using an induction on $n$ that exploits the Heisenberg fiber structure together with a multilinear interpolation scheme. Specializing to indicator functions yields a sharp Loomis–Whitney type set inequality bounding $|K|$ for every finite $K\subset\mathbb{H}^{n}(\mathbb{F}_{q})$ in terms of the sizes of its $2n$ Heisenberg projections $\{\pi_{j}(K)\}_{j=1}^{2n}$, and in particular,
$\max_{1\leq j\leq 2n}|\pi_{j}(K)|\;\gtrsim_{n}\;|K|^{\frac{2n+1}{2(n+1)}}\,q^{-\frac{1}{2(n+1)}}.$
This result is optimal up to absolute constants. Moreover, when $n=1$ and $|K|>q$, we obtain a stronger statement via Vinh’s point–line incidence theorem. We also discuss connections to a boundedness problem for multilinear forms/operators over finite fields studied by Bhowmik, Iosevich, Koh, and Pham (2025), and to orthogonal projection/covering questions in $\mathbb{F}_{q}^{2n+1}$ studied by Chen (2018).
@@ 1 Introduction
The Loomis–Whitney inequality ([15]) bounds the volume of a dd–dimensional body in $\mathbb{R}^{d}$ by the volumes of its (d1)(d-1)–dimensional coordinate projections. In the discrete setting, it reads as follows. For $d\geq 2$, let $B_{1},\dots,B_{d}$ be finite sets, and let $A\subseteq B_{1}\times\cdots\times B_{d}$. For $i\in\{1,\ldots,d\}$, define
@@
There are powerful entropy-theoretic generalizations of (1), notably Han’s inequality and Shearer’s inequality [9, 16]. Variants of this inequality, in relation to set addition, also appear in additive combinatorics for abelian groups with numerous applications [8, 18].
This paper investigates Loomis–Whitney type inequalities in the noncommutative setting, focusing on the Heisenberg group over a finite field $\mathbb{F}_{q}$, where $q$ is an odd prime power. In this context, such inequalities are not only of independent interest but also shed light on group structures and are connected to structural properties of the Heisenberg group and to related combinatorial problems. Related works on the Heisenberg group can be found in [1, 11, 12, 17]. Interestingly, this topic is naturally connected to boundedness problems for multilinear forms/operators over finite fields and to orthogonal projection/covering questions in $\mathbb{F}_{q}^{2n+1}$. Further details will be discussed in Section 7.
In the continuous setting, the Loomis–Whitney type inequalities for the Heisenberg group (and, more generally, Carnot group) have been studied extensively with applications in multilinear restriction theory and related areas; for example, see [3, 7, 21] and references therein.
@@
By definition, the group $\mathbb{H}^{n}(\mathbb{F}_{q})$ and the vector space $\mathbb{F}_{q}^{2n+1}$ have the same underlying set. For each $1\leq j\leq 2n$, let
@@
be the coordinate line corresponding to $x_{j}$, where $\{e_{j}\}_{j=1}^{2n}$ form the standard basis of $\mathbb{F}_{q}^{2n}$.
Then we easily see that $W_{j}$ and $L_{j}$ are subgroups of $\mathbb{H}^{n}(\mathbb{F}_{q})$ for $j\in\{1,\ldots,2n\}$. Further, $L_{j}$ is the orthogonal complement $W_{j}^{\perp}$ of $W_{j}$, with respect to the standard inner product on $\mathbb{F}_{q}^{2n+1}$; see Subsection 2.2 for these materials. The subgroups $W_{j}$ and $L_{j}$ will be often referred to as vertical and horizontal, respectively. We will provide more details on subgroup structures of $\mathbb{H}^{n}(\mathbb{F}_{q})$ in Section 2.
For $x\in\mathbb{F}^{2n}_{q}$ and $j\in\{1,\ldots,2n\}$, we use the notation $\widehat{x}_{j}$ to represent the element in $\mathbb{F}^{2n-1}_{q}$ obtained by entirely removing the $j$-th coordinate of $x$ while the notation $\check{x}_{j}$ represents the element in $\mathbb{F}^{2n}_{q}$ obtained by setting the $j$-th coordinate of $x$ to 0. For $j\in\{1,\ldots,n\}$, let
@@
be the vertical coordinate projections, respectively, defined by
@@
A direct computation shows that for each $j\in\{1,\ldots,2n\}$, an element $(x,t)$ of $\mathbb{H}^{n}(\mathbb{F}_{q})$ can be written
@@
The fiber of $\pi_{j}$ at a point $(u,\tau)$ in $W_{j}$ is a line in the affine space $\mathbb{F}_{q}^{2n+1}$ passing through $(u,\tau)$. More precisely, for $j\in\{1,\ldots,n\}$, the fiber of $\pi_{j}$ at $(u,\tau)\in W_{j}$ (so $u_{j}=0$) is given by
@@
That is, for each $j\in\{1,\ldots,2n\}$, the fibers of $\pi_{j}$ are exactly cosets of $W_{j}^{\perp}=L_{j}$.
@@
Throughout, for each $j\in\{1,\ldots,2n\}$, we shall identify $W_{j}$ with $\mathbb{F}_{q}^{2n}$ via the map $\phi_{j}:W_{j}\stackrel{{\scriptstyle\sim}}{{\rightarrow}}\mathbb{F}_{q}^{2n}$ defined by $(\check{x}_{j},t)\mapsto(\widehat{x}_{j},t)$. For simplicity, for a function $f$ on $\mathbb{F}_{q}^{2n}$ and $(x,t)\in\mathbb{H}^{n}(\mathbb{F}_{q})$, we write $f(\pi_{j}(x,t))$ for $f(\phi_{j}(\pi_{j}(x,t)))$.
@@
Let $n$ be a positive integer. For extended real numbers $1\leq u_{1},u_{2},\ldots,u_{2n}\leq\infty$, define $LW(u_{1},u_{2},\ldots,u_{2n})$ as the smallest constant not depending on $q$, if it exists, such that the inequality
@@
holds for all nonnegative $f_{1},f_{2},\ldots,f_{2n}\in\mathscr{F}(\mathbb{F}_{q}^{2n})$. If it does not exist, then we set $LW(u_{1},u_{2},\ldots,u_{2n})=\infty.$
Although our main results are stated as multilinear functional inequalities, they immediately imply corresponding set/projection statements. When $f_{1}=\cdots=f_{2n}$ are all the indicator functions of a subset $K$ of $\mathbb{H}^{n}(\mathbb{F}_{q})$, then the inequality (7) gives a Loomis–Whitney type inequality.
@@
In this paper, we write $X\lesssim Y$ to mean that there exists some constant $C>0$ such that $X\leq CY$.
@@
For any nonnegative $f_{1},f_{2}\in\mathscr{F}(\mathbb{F}_{q}^{2})$,
$\frac{1}{q^{3}}\sum_{(x,t)\in\mathbb{H}^{1}(\mathbb{F}_{q})}f_{1}(\pi_{1}(x,t))\,f_{2}(\pi_{2}(x,t))\;\lesssim\;\|f_{1}\|_{L^{\frac{3}{2}}(\mathbb{F}_{q}^{2},dx)}\,\|f_{2}\|_{L^{\frac{3}{2}}(\mathbb{F}_{q}^{2},dx)}.$
The endpoint exponent $\frac{3}{2}$ is optimal for general functions $f_{1},f_{2}$ by Theorem 1.2. On the level of sets, Corollary 1.3 implies the bound
$|K|\lesssim q^{\frac{1}{3}}\,|\pi_{1}(K)|^{\frac{2}{3}}\,|\pi_{2}(K)|^{\frac{2}{3}},\qquad K\subset\mathbb{H}^{1}(\mathbb{F}_{q}),$ (8)
and this is sharp up to absolute constants. Indeed, fixing $t_{0}\in\mathbb{F}_{q}$ and taking
$K:=\{(x_{1},0,t_{0})\in\mathbb{H}^{1}(\mathbb{F}_{q})\colon x_{1}\in\mathbb{F}_{q}\},$
we have $|K|=q$, $|\pi_{1}(K)|=1$, and $|\pi_{2}(K)|=q$, so (8) is attained.
This construction can be extended to higher dimensions as follows. Let
$K:=\{(x,t_{0})\in\mathbb{H}^{n}(\mathbb{F}_{q}):x_{n+1}=\cdots=x_{2n}=0\}.$
Then, $|K|=q^{n}$, and
This example naturally suggests the critical symmetric exponent in higher dimensions. Suppose one seeks a uniform estimate of the form
$\frac{1}{q^{2n+1}}\sum_{(x,t)\in\mathbb{H}^{n}(\mathbb{F}_{q})}\prod_{j=1}^{2n}f_{j}(\pi_{j}(x,t))\;\lesssim\;\prod_{j=1}^{2n}\|f_{j}\|_{L^{u}(\mathbb{F}_{q}^{2n},dx)}$ (9)
with a single exponent $u$ for all inputs. Applying (9) to indicator functions $f_{j}=\mathbf{1}_{\pi_{j}(K)}$ gives
$\frac{|K|}{q^{2n+1}}\;\lesssim\;\prod_{j=1}^{2n}\Bigg(\frac{|\pi_{j}(K)|}{q^{2n}}\Bigg)^{1/u}.$
Inserting the above values of $|K|$ and $|\pi_{j}(K)|$ forces
$q^{n}\;\lesssim\;q^{\,2n+1-\frac{2n^{2}+n}{u}},\qquad\text{hence}\qquad u\;\geq\;\frac{n(2n+1)}{n+1}.$
Thus, $\frac{n(2n+1)}{n+1}$ is the smallest possible symmetric exponent one can expect when $n>1$.
When $n>1$, the full description of the admissible region of $(u_{1},\dots,u_{2n})$ for which $LW(u_{1},\dots,u_{2n})\lesssim 1$ holds is substantially more intricate than in the planar case $n=1$. Rather than attempting a complete characterization, in this paper, we focus on the symmetric critical value $u=\frac{n(2n+1)}{n+1}$, for which the inequality is established in the next theorem.
@@
The proof of the corresponding 2n2n–multilinear inequality is not a formal iteration of the case $n=1$. It is inspired by the induction (on dimension) scheme developed in the continuous setting [7], but requires nontrivial adaptations to the finite field setting and to obtaining constants uniform in $q$. It has two main components. First, we establish a family of asymmetric estimates (Theorem 4.1) in which one symplectic pair of inputs is measured in $L^{\frac{2n+1}{2}}$ while the remaining inputs are measured in $L^{2n+1}$. These bounds are proved by a delicate induction on $n$ that repeatedly freezes one symplectic coordinate pair and rewrites the nn–dimensional form as an averaged (n1)(n-1)–dimensional Heisenberg Loomis–Whitney form. Second, we combine the resulting family of asymmetric bounds through a multilinear interpolation scheme to reach the symmetric estimate at the critical exponent $u=\frac{n(2n+1)}{n+1}$.
@@
It follows from the example above that this corollary is sharp and we cannot replace the factor $q^{\frac{1}{2n+1}}$ by $|\pi_{v}(K)|^{\frac{1}{2n+1}}$, where $\pi_{v}(K)$ is the image of $K$ under the (vertical) projection $\pi_{v}:\mathbb{H}^{n}(\mathbb{F}_{q})\rightarrow\mathbb{F}_{q}$ defined by $(x,t)\mapsto t$.
When $n=1$ and $|K|>q$, by using the point–line incidence theorem due to Vinh [20], we are able to provide a better lower bound for the quantity $\max\{|\pi_{1}(K)|,|\pi_{2}(K)|\}$.
@@
$\frac{|\pi_{1}(K)|\,|\pi_{2}(K)|}{q}\;+\;q^{\frac{1}{2}}\sqrt{|\pi_{1}(K)|\,|\pi_{2}(K)|}\;=\;m^{2}q\;+\;mq^{\frac{3}{2}},$